OpenStudy (anonymous):

What is the original length of a 4x6 note card?

OpenStudy (anonymous):

huh?

OpenStudy (wolf1728):

Ummm 4 by 6?

OpenStudy (anonymous):

Well these are the question i need help with

OpenStudy (wolf1728):

Wow - sorry for making a little fun there. Oh well I see that comes with an explanation and a diagram. That changes things considerably. Original length is 6" (or 6 inches).

OpenStudy (anonymous):

lol its okay, i know im not smart in any type of math /:

OpenStudy (anonymous):

so then the original length is 6'' and the original width is 4" ?

OpenStudy (wolf1728):

B) Four pieces have been cut. C) We are not told the length of 'x' so we have to call the resulting length 6 inches minus 2x. D) Original width - 4 inches. E) Four pieces have been cut. F) Resulting width = 4 inches minus 2x.

OpenStudy (anonymous):

thank you so much! for C, would the resulting length be 2 since it was originally 6 inches and four pieces were cut?

OpenStudy (wolf1728):

Since we don't know the length of 'x', we can't say definitely what the length or width is. If x is a half inch the resulting width = 6 minus 2*half = 6 -1 = 5.

OpenStudy (anonymous):

oh, i see, that completely makes sense! thank you so much! :)

OpenStudy (wolf1728):

Okay, glad to have helped out. :-)

OpenStudy (anonymous):

well now I am confused because the next question says "Use your answers to parts c and f to express the new dimensions of the box. "

OpenStudy (anonymous):

how do i do that if we dont know the true length and width?

OpenStudy (wolf1728):

I'd answer it by what I stated in part c and f

OpenStudy (anonymous):

so then it would 6''-2x by 4"-2x?

OpenStudy (wolf1728):

Yes or as it might be stated in algebra (6-2x)" by (4-2x)"

OpenStudy (anonymous):

that makes sense, but the last question in that section is "The area of the bottom of the box is the product of the new length and width. Write this area as a function in the form: A(x) = your answer"

OpenStudy (anonymous):

so how would i go about that if i dont know the exact dimensions of the box?

OpenStudy (wolf1728):

Multiply (6-2x)" by (4-2x)" Is this an intro to algebra course?

OpenStudy (anonymous):

it's pre calculus

OpenStudy (wolf1728):

I guess they want you to think in the abstract. Let me multiply (6-2x)" by (4-2x)"

OpenStudy (anonymous):

okay!

OpenStudy (wolf1728):

If we say x is 3/4 of an inch then width = 4 - 2*.75 = 2.5" and length = 6 -2*.75 = 6 -1.5 = 4.5" The area of the bottom of that box is 2.5 times 4.5 * 2.5 = 11.25

OpenStudy (wolf1728):

(6-2x)" by (4-2x)" = 24 -8x -12x -4x² = 24 -20x +4x² Let's say x = .75 24 -20*.75 +4*.75*.75 =3.75 + 2.25 =6

OpenStudy (anonymous):

thank you so much!! that was very helpful :)

OpenStudy (wolf1728):

(6-2x)" by (4-2x)" = 24 -8x -12x -4x² = 24 -20x +4x² Let's say x = .75 24 -20*.75 +4*.75*.75 = 24 -15 + 2.25 = 11.25

OpenStudy (anonymous):

Could you help me with these 2 questions? I promise these are the last ones for this problem! lol it is just so confusing

OpenStudy (wolf1728):

It seems there must be a Part II to that attachment because it talks about "set your question from Part II"

OpenStudy (anonymous):

we just did part II

OpenStudy (wolf1728):

No - that was Part I

OpenStudy (anonymous):

Part I-III

OpenStudy (wolf1728):

See it says "Part I (6 points)"

OpenStudy (anonymous):

Yes, and then right under it is part II for 8 points

OpenStudy (wolf1728):

My answer (I don't think) was submitted by you and I think it can be used for part II A and B (6-2x)" by (4-2x)" = 24 -8x -12x +4x² = 24 -20x +4x²

OpenStudy (wolf1728):

Part III A. and B. Setting this up in conventional form: Area = 4x² -20x + 24 If we want an area of 16 square inches then 16 = 4x² -20x + 24 4x² -20x + 8 Solving by quadratic formula a=4 b=-20 c=8 x = [--20 +-sqrt(400 - 4 * 4 * 8)] / 2*4 x = [20 +- sqrt(272)] / 8 The 2 answers are: x1 = 4.5615528128 and x2 = 0.4384471872 Answer x1 can't be used because 2 * 4.56 is greater than 4 or 6 so the note card couldn't have ANY dimension. So the only usable answer is x2 = 0.4384471872

OpenStudy (anonymous):

thank you so much!

OpenStudy (wolf1728):

u r welcome and thanks for the medal :-)